The sum of money shared among nine people so that the first gets 75, 000, the next 150, 000, the next 225, 000, and so on. (a) How much money does the ninth person get? (b) How much money is shared altogether?
Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



The amount the n-th person gets is 75,000 times n.

The amount the 9th person gets is 9*75,000 = 675,000.

The total amount is 75,000(1+2+...+9) = 75,000(45) = 3,375,000.

Answer by ikleyn(52788)

.
The amount the n-th person gets is 75,000 times n.

The amount the 9th person gets is 9*75,000 = 675,000.              ANSWER, (a)

The total amount is 75,000(1+2+...+9) = 75,000(45) = 3,375,000.    ANSWER, (b)